OpenStudy (unklerhaukus):

\[y^{\prime\prime}-3y^\prime+2y=e^x\]

OpenStudy (unklerhaukus):

\[y_c^{\prime\prime}-3y_c^\prime+2y_c=0\] \[m^2-3m+2=0\]\[(m-1)(m-2)=0\]\[m=1,2\] \[y_c=Ae^x+Be^{2x}\]

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}-3y_p^\prime+2y_p=e^x\] \[y_p=Cxe^x\]\[y^\prime_p=C(1+x)e^x\]\[y^{\prime\prime}_p=C(2+x)e^x\] \[C(2+x)e^x-3C(1+x)e^x+2Cxe^x=e^x\]\[C\left(2+x-3-3x+2x\right)e^{x}=e^x\]\[-C=1;\qquad C=-1\] \[y_p=-xe^x\]

OpenStudy (unklerhaukus):

\[y=y_c+y_p\] \[y(x)=Ae^x+Be^{2x}-xe^x\]

OpenStudy (unklerhaukus):

my text book has this as the solution: \[y(x)=Ae^x+Be^{2x}-(x+1)e^x\]

OpenStudy (unklerhaukus):

is the book right/? where doe the extra one come from?

OpenStudy (anonymous):

I think you cannot take particular solution as \(e^x\) because it is in the homogeneous solution too..

OpenStudy (anonymous):

But no, yes you can take because you are multiplying x with it..

OpenStudy (unklerhaukus):

i took the particular as \(y_p=Cxe^x\)

OpenStudy (turingtest):

\[A _1e^x-e^x=A_2e^x\]so the constant for the other e^x part gets "absorbed" into the other

OpenStudy (turingtest):

no need to write both; they are linearly dependent

OpenStudy (valpey):

They are both correct.

OpenStudy (anonymous):

Yes, I got that Uncle Rocks.. I tried by using \((Cx + D)e^x\) then I got \(C = -1 \quad and \quad D = 0\) then also.. You are right @UnkleRhaukus

OpenStudy (turingtest):

\[y(x)=Ae^x+Be^{2x}-xe^x-e^x=(A-1)e^x+Be^x=Ce^x+Be^{2x}\]same thing, yup :)

OpenStudy (turingtest):

\[y(x)=Ae^x+Be^{2x}-xe^x-e^x\]\[=(A-1)e^x+Be^x-xe^x\]\[=Ce^x+Be^{2x}-xe^x\]

OpenStudy (anonymous):

Again typo in middle line... Ha ha ha..

OpenStudy (turingtest):

yes there is :S

OpenStudy (unklerhaukus):

forgot a 2

OpenStudy (turingtest):

but I hope you can fill in the missing 2

OpenStudy (turingtest):

it's 1:16 am here folks, gimme a break :P

OpenStudy (unklerhaukus):

i think i get it now, the book is just using a constant that is different to my constant by one unit,

OpenStudy (turingtest):

yeah that's for sure the deal I just encountered this same coefficient problem recently, so I was able to spot it pretty quickly

OpenStudy (unklerhaukus):

so, i'll just leave the solution how i had it \[y(x)=Ae^x+Be^{2x}-xe^x\] is neater than \[y(x)=A'e^x+Be^{2x}-(x+1)e^x\]

OpenStudy (valpey):

Would the book be advocating?: \[y_p=Cxe^x+e^x\]

OpenStudy (unklerhaukus):

i am not sure @Valpey

OpenStudy (anonymous):

@UnkleRhaukus turning test is saying right.. You have done it correctly..

OpenStudy (unklerhaukus):

im just not sure how the book go to what it did, or why i took that approach

OpenStudy (unklerhaukus):

why it*

OpenStudy (anonymous):

p=1,m=0 Unkle,Do you remember?

OpenStudy (unklerhaukus):

i do not remember

OpenStudy (anonymous):

take a look the post that help you for all problems. Amazing individual way.

OpenStudy (unklerhaukus):

what do yon mean by p and m

OpenStudy (anonymous):